Standing waves are established in a string. The frequency of the wave is 512 Hz and the 

distance between two successive nodes is 8 cm. Calculate the velocity of the wave in the 

string.



1
Expert's answer
2021-05-17T19:01:43-0400

velocity of wave = frequency "\times" wavelength



wavelength = 8 cm = 0.08 m


Frequency(n)= 512 Hz




the distance between two nodes represents half wavelength = "\frac{1}{2}\lambda"

v="n\lambda"



wavelength = 2"\times" 0.08 = 0.16 "m"


v="n\lambda"

v="512\times0.16" m/sec

v=81.92m/sec

Velocity of wave

v=81.92m/s